# Minimax rates of convergence for nonparametric regression under adversarial attacks

### v1

@@ A.1.4 Proof of Theorem 3
$R_{\Delta_{n}}(f,\widehat{f})\geq L_{\Delta_{n}}^{*}(f).$ (A.1.6)
In addition, the adversarial risk is always lower bounded by the standard risk, i.e.,
@@
Appendix A.2 Proofs of the results in two examples
A.2.1 Proof of Example 1
@@
denote its Taylor polynomial of degree $k=\lfloor\beta\rfloor$ of at point $t$. Using results from the approximation theory (see, e.g., DeVore and Lorentz,, 1993), we know that
@@
$\begin{split}2L_{\Delta_{n}}^{*}(f)&=\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right].\\ \end{split}$
@@
$\begin{split}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|f(x)-g_{k}(x^{\prime};\bar{x})\right|&\leq\frac{L}{2^{\beta}}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{|l|=k}\frac{1}{l!}\left|x-x^{\prime}\right|^{l}\cdot\left\|x-x^{\prime}\right\|^{\alpha}\\ &\leq\frac{Lr_{n}^{\alpha}}{2^{k}}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{|l|=k}\frac{1}{l!}\left|x-x^{\prime}\right|^{l}\\ &=\frac{Lr_{n}^{\alpha}}{2^{k}k!}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left(|x_{1}-x^{\prime}_{1}|+\cdots+|x_{d}-x^{\prime}_{d}|\right)^{k}\\ &\leq\frac{Lr_{n}^{\alpha}d^{\frac{k}{2}}}{2^{k}k!}\left(|x_{1}-x^{\prime}_{1}|^{2}+\cdots+|x_{d}-x^{\prime}_{d}|^{2}\right)^{\frac{k}{2}}\\ &\leq\frac{Ld^{\frac{k}{2}}r_{n}^{\alpha}(2r_{n})^{k}}{2^{k}k!}\leq Cr_{n}^{\beta},\end{split}$ (A.2.4)
@@
$\begin{split}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|g_{k}(x;\bar{x})-g_{k}(x^{\prime};\bar{x})\right|&=\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|\sum_{|l|\leq k}\frac{D^{l}f(\bar{x})}{l!}\left[(x-\bar{x})^{l}-(x^{\prime}-\bar{x})^{l}\right]\right|\\ &=\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|\sum_{|l|\leq k}\frac{D^{l}f(\bar{x})}{2^{|l|}l!}\left[1+(-1)^{|l|+1}\right](x-x^{\prime})^{l}\right|\\ &\leq C\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{1\leq s\leq k}\sum_{|l|=s}\frac{1}{l!}|x-x^{\prime}|^{l}\\ &\leq C\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{1\leq s\leq k}\left(|x_{1}-x^{\prime}_{1}|^{2}+\cdots+|x_{d}-x^{\prime}_{d}|^{2}\right)^{\frac{s}{2}}\\ &\leq Cr_{n},\end{split}$ (A.2.5)
@@
$\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|g_{k}(x^{\prime};\bar{x})-f(x^{\prime})\right|\leq Cr_{n}^{\beta}.$ (A.2.6)
Combining (A.2.3) with (A.2.4)–(A.2.6), we have for all $f\in\mathcal{F}_{1}(\beta,L)$, $L_{\Delta_{n}}^{*}(f)\leq Cr_{n}^{1\wedge\beta}$, i.e.,
$\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)\leq Cr_{n}^{1\wedge\beta}.$
To lower bound $\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)$, it suffices to construct specific functions in $\mathcal{F}_{1}(\beta,L)$ such that $L_{\Delta_{n}}^{*}(f)\geq Cr_{n}^{1\wedge\beta}$. Given that $\Delta_{n}$ is a closed set, let $\delta$ and $\delta^{\prime}$ be any two points in $\Delta_{n}$ such that $\|\delta-\delta^{\prime}\|=r_{n}$. In addition, define $D_{n}=\{t\delta+(1-t)\delta^{\prime}:0\leq t\leq 1\}$. Since there exists a $D_{n}$ such that $D_{n}\subseteq\Delta_{n}$, hence $L_{\Delta_{n}}^{*}(f)\geq L_{D_{n}}^{*}(f)$. Without loss of generality, we assume $D_{n}\subseteq\{x:x_{2}=x_{3}=\cdots=x_{d}=0\}$ and $(\delta+\delta^{\prime})/2=(1/2,0,\ldots,0)^{\top}$. Otherwise, we can construct new functions from the functions $f_{1}$ and $f_{2}$ defined below by rotations of axes. Note that the rotation and transformation of a function do not change the smoothness properties of the original function. When $\beta\geq 1$, define
@@
$\begin{split}\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)\geq L_{D_{n}}^{*}(f_{1})&\geq C\left[\sup_{x\in\{(1/2,0,\ldots,0)^{\top}-D_{n}\}}f_{1}(x)-\inf_{x\in\{(1/2,0,\ldots,0)^{\top}-D_{n}\}}f_{1}(x)\right]\\ &\geq Cr_{n}.\end{split}$
@@ A.2.2 Proof of Example 2
$\begin{split}2L_{\Delta_{n}}^{*}(f)&=\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right]\\ &\leq\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x,z\in x^{\prime}-\Delta_{n}}\left|f(x)-f(z)\right|\right]\\ &\leq\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x,z\in x^{\prime}-\Delta_{n}}\left(L_{1}\left|x_{1}-z_{1}\right|^{\beta_{1}}+\cdots+L_{d}\left|x_{d}-z_{d}\right|^{\beta_{d}}\right)\right]\\ &\leq L_{1}r_{1}^{\beta_{1}}+\cdots+L_{d}r_{d}^{\beta_{d}}\\ &\lesssim\max\{r_{1}^{\beta_{1}},\ldots,r_{d}^{\beta_{d}}\},\end{split}$ (A.2.8)

### v2

@@ Appendix A.4 Proof of Theorem 3
$\begin{split}R_{\Delta_{n}}(f,\widehat{f})&\geq\sup_{x^{\prime}\in\mathcal{X}}\left[\frac{\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)}{2}\right]\\ &=L_{\Delta_{n}}^{*}(f),\end{split}$ (A.4.1)
where the equality follows from (3.2). In addition, the adversarial risk is always lower bounded by the standard risk, i.e.,
@@
Appendix A.5 Proof of Example 1
@@
denote its Taylor polynomial of degree $k=\lfloor\beta\rfloor$ at point $t$. Using results from the approximation theory (see, e.g., DeVore and Lorentz,, 1993), we know that
@@
$\begin{split}2L_{\Delta_{n}}^{*}(f)&=\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right].\\ \end{split}$
@@
$\begin{split}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|f(x)-g_{k}(x^{\prime};\bar{x})\right|&\leq\frac{L}{2^{\beta}}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{|l|=k}\frac{1}{l!}\left|x-x^{\prime}\right|^{l}\cdot\left\|x-x^{\prime}\right\|^{\alpha}\\ &\leq\frac{Lr_{n}^{\alpha}}{2^{k}}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{|l|=k}\frac{1}{l!}\left|x-x^{\prime}\right|^{l}\\ &=\frac{Lr_{n}^{\alpha}}{2^{k}k!}\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left(|x_{1}-x^{\prime}_{1}|+\cdots+|x_{d}-x^{\prime}_{d}|\right)^{k}\\ &\leq\frac{Lr_{n}^{\alpha}d^{\frac{k}{2}}}{2^{k}k!}\left(|x_{1}-x^{\prime}_{1}|^{2}+\cdots+|x_{d}-x^{\prime}_{d}|^{2}\right)^{\frac{k}{2}}\\ &\leq\frac{Ld^{\frac{k}{2}}r_{n}^{\alpha}(2r_{n})^{k}}{2^{k}k!}\leq Cd^{\frac{k}{2}}r_{n}^{\beta},\end{split}$ (A.5.4)
@@
$\begin{split}&\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|g_{k}(x;\bar{x})-g_{k}(x^{\prime};\bar{x})\right|\\ &=\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|\sum_{|l|\leq k}\frac{D^{l}f(\bar{x})}{l!}\left[(x-\bar{x})^{l}-(x^{\prime}-\bar{x})^{l}\right]\right|\\ &=\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|\sum_{|l|\leq k}\frac{D^{l}f(\bar{x})}{2^{|l|}l!}\left[1+(-1)^{|l|+1}\right](x-x^{\prime})^{l}\right|\\ &\leq C\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{1\leq s\leq k}\sum_{|l|=s}\frac{1}{l!}|x-x^{\prime}|^{l}\\ &\leq C\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\sum_{1\leq s\leq k}\left(|x_{1}-x^{\prime}_{1}|^{2}+\cdots+|x_{d}-x^{\prime}_{d}|^{2}\right)^{\frac{s}{2}}\\ &\leq Cr_{n},\end{split}$ (A.5.5)
@@
$\sup_{\|x-x^{\prime}\|\leq 2r_{n}}\left|g_{k}(x^{\prime};\bar{x})-f(x^{\prime})\right|\leq Cd^{\frac{k}{2}}r_{n}^{\beta}.$ (A.5.6)
Combining (A.5.3) with (A.5.4)–(A.5.6), we have for all $f\in\mathcal{F}_{1}(\beta,L)$, $L_{\Delta_{n}}^{*}(f)\leq Cd^{\frac{k}{2}}r_{n}^{1\wedge\beta}$, i.e.,
$\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)\leq Cd^{\frac{k}{2}}r_{n}^{1\wedge\beta}.$
To lower bound $\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)$, it suffices to construct specific functions in $\mathcal{F}_{1}(\beta,L)$ such that $L_{\Delta_{n}}^{*}(f)\geq Cr_{n}^{1\wedge\beta}$. Given that $\Delta_{n}$ is a closed set, let $\delta$ and $\delta^{\prime}$ be two points in $\Delta_{n}$ such that $\|\delta-\delta^{\prime}\|=r_{n}$. In addition, define $D_{n}=\{t\delta+(1-t)\delta^{\prime}:0\leq t\leq 1\}$. Since there exists a $D_{n}$ such that $D_{n}\subseteq\Delta_{n}$, hence $L_{\Delta_{n}}^{*}(f)\geq L_{D_{n}}^{*}(f)$. Without loss of generality, we assume $D_{n}\subseteq\{x:x_{2}=x_{3}=\cdots=x_{d}=0\}$ and $(\delta+\delta^{\prime})/2=(1/2,0,\ldots,0)^{\top}$. Otherwise, we can construct new functions from the functions $f_{1}$ and $f_{2}$ defined below by rotations of axes and shifts of origin. Note that the rotation and transformation of a function do not change the smoothness properties of the original function. When $\beta\geq 1$, define
@@
$\begin{split}\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)\geq L_{D_{n}}^{*}(f_{1})&\geq C\left[\sup_{x\in\{(1/2,0,\ldots,0)^{\top}-D_{n}\}}f_{1}(x)\right.\\ &\qquad\qquad\qquad\left.-\inf_{x\in\{(1/2,0,\ldots,0)^{\top}-D_{n}\}}f_{1}(x)\right]\\ &\geq Cr_{n}.\end{split}$
@@ Appendix A.6 Proof of Example 2
$\begin{split}2L_{\Delta_{n}}^{*}(f)&=\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right]\\ &\leq\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x,z\in x^{\prime}-\Delta_{n}}\left|f(x)-f(z)\right|\right]\\ &\leq\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x,z\in x^{\prime}-\Delta_{n}}\left(L_{1}\left|x_{1}-z_{1}\right|^{\beta_{1}}+\cdots+L_{d}\left|x_{d}-z_{d}\right|^{\beta_{d}}\right)\right]\\ &\leq L_{1}r_{1}^{\beta_{1}}+\cdots+L_{d}r_{d}^{\beta_{d}}\\ &\lesssim\max\{r_{1}^{\beta_{1}},\ldots,r_{d}^{\beta_{d}}\},\end{split}$ (A.6.2)